# Improving Human Verification of LLM Reasoning through Interactive Explanation Interfaces

### v3

@@ Abstract
The reasoning capabilities of Large Language Models (LLMs) have led to their increasing employment in several critical applications, particularly education, where they support problem-solving, tutoring, and personalized study. Chain-of-thought (CoT) reasoning capabilities [guo2025deepseek, wei2022chain] are well-known to help LLMs decompose a problem into steps and explore the solution spaces more effectively, leading to impressive performance on mathematical and reasoning benchmarks. As the length of CoT tokens per question increases substantially to even thousands of tokens per question [guo2025deepseek], it is unknown how users could comprehend LLM reasoning and detect errors or hallucinations. To address this problem and understand how reasoning can improve human-AI interaction, we present three new interactive reasoning interfaces: interactive CoT (iCoT), interactive Program-of-Thought (iPoT), and interactive Graph (iGraph). That is, we ask LLMs themselves to generate an interactive web interface wrapped around the original CoT content, which may be presented in text (iCoT), graphs (iGraph) or code (iPoT). This interface allows users to interact with and provide a novel experience in reading and validating the reasoning chains of LLMs. Across a study of 125 participants, interactive interfaces significantly improve user performance. Specifically, iGraph users score the highest error detection rate ($85.6\%$), followed by iPoT ($82.5\%$), iCoT ($80.6\%$), all outperforming standard CoT ($73.5\%$). Interactive interfaces also lead to faster user validation time—iGraph users are faster ($57.9$ secs per question) than the users of iCoT and iPoT ($60$ secs) and the standard CoT ($64.7$ secs). A post-study questionnaire shows that users prefer iGraph, citing its superior ability to enable them to follow the LLM’s reasoning. We discuss the implications of these results and provide recommendations for the future design of reasoning models. Try our interactive interfaces on HuggingFace.
1 Introduction
@@
Large language models (LLMs) are now woven into high-stakes, real-world workflows, including education, where they assist with problem solving, tutoring, and personalized study [kojima2022large, wang2024large, forootani2025survey]. A key driver of their recent success on complex tasks is chain-of-thought (CoT) prompting [wang-etal-2023-towards]: eliciting step-by-step intermediate reasoning often yields dramatic accuracy gains on arithmetic, commonsense, and symbolic reasoning benchmarks, with state-of-the-art performance on GSM8K and related math datasets. However, the current format for presenting LLM explanations may be far from optimal [agarwal2024faithfulness].
@@
Figure 1: Given a GSM8K question, LLMs typically provide step-by-step reasoning followed by the final answer. However, such output presentation is often static and long, posing a higher cognitive load to users and leading to slower and more erroneous answer verification. In contrast, we prompt LLMs to generate an interactive HTML/JavaScript application wrapped around the reasoning. This interface enables users to verify the reasoning more efficiently via tools of (a) navigation buttons (inspired by those in common IDEs) and (b) colored highlights [nguyen2025hot].
@@
While CoT outputs contain rich information, their long-text presentation introduces significant challenges for human interpretation [kunz2024properties, hedlin2025got, turpin2023language]. In practice, especially in modern reasoning LLMs [guo2025deepseek], CoT often appears as a long “wall of text” that can be time-consuming, obscure structure, increase cognitive load, and make error detection harder precisely when it matters. For example, when an explanation contains a subtle mistake in one intermediate step, detecting that error often requires carefully parsing multiple sentences and mentally reconstructing the reasoning flow [mishra2025investigating]. More broadly, the explainable AI (XAI) literature shows that explanation techniques are frequently designed around what researchers can produce rather than what end users need; users instead report wanting practically useful information that improves collaboration, trust calibration, and feedback loops with AI systems [miller2019explanation, abdul2018trends, kaur2020interpreting]. This gap motivates rethinking how reasoning is surfaced to end-users in contexts like education, where clarity, verification, and engagement are essential.
@@
Prior work has demonstrated that the design of LLM explanations, which includes factors such as visual structure, interactivity, and information hierarchy, can significantly influence users’ trust and understanding [yang2020visual, alicioglu2022survey, jiang2023graphologue, nguyen2024interpretable]. However, most existing research on CoT focuses on improving the reasoning process within the model rather than optimizing the presentation of the reasoning [wang2024chain, wang2022towards]. As a result, LLM explanations remain largely static, text-heavy, and difficult to navigate [zhao2024explainability]. This disconnect raises critical questions: How can explanation formats be redesigned to reduce cognitive load and enable users to effectively detect errors and understand LLM-generated reasoning steps?
@@
RQ1) Do interactive explanation formats help users improve verification accuracy compared to traditional text outputs?
RQ2) How do these formats influence the time and effort required for users to verify reasoning?
@@ 2 Related work
LLMs in Education with Chain-of-Thought (CoT) Prompting. A key prompting technique for LLMs is CoT prompting, which elicits the LLM’s reasoning steps and improves its performance on multi-step tasks [zhang2022automatic]. A recent study shows that few-shot CoT dramatically boosts arithmetic, commonsense, and symbolic reasoning accuracy [wei2022chain, kojima2022large]. In educational settings, however, how explanations are presented matters. Despite clear opportunities for tutoring and feedback, khosravi2022explainable highlights persistent concerns about transparency, interpretability, and over-reliance in LLMs. Beyond transparency, CoT reasoning tends to be quite lengthy, which can increase cognitive load and slow down the user’s comprehension [nayab2024concise]. Given the concerns about the interpretability and cognitive demands of LLM-generated reasoning, researchers have begun to explore how reasoning quality and user understanding can be evaluated more systematically. A growing body of work has adopted verification accuracy, which indicates the users’ ability to correctly judge whether a model’s reasoning is valid, as a key indicator of comprehension in LLM explanations [narayanan2018humans, linder2021level, nourani2020don, lage2019human, nguyen2024interpretable]. Further, a recent study argues that explanations are only useful when they enable users to verify predictions step-by-step, rather than simply accepting plausible reasoning at face value [fok2024search]. In our study, we aim to address this question by measuring how effectively people can detect reasoning errors across different explanation formats, allowing us to assess not only the quality of the reasoning itself but also how various interface designs influence users’ ability to verify and understand it.
@@
Interactivity in LLM Explanations. Recent studies have emphasized the importance of interactivity in AI explanations to improve deeper user engagement and understanding[kim2020improved, guesmi2024interactive, nguyen2024allowing]teso2023leveraging provides a comprehensive overview of interactive machine learning, which shows that explanations are most effective when users can query, edit, or refine them rather than passively receive them. They argue that interaction transforms explanations from static outputs into dynamic learning tools that encourage users’ cognitive engagement. Similarly, another work reviews human-centered explainability and finds that interactive explanation interfaces significantly enhance engagement, comprehension, and retention compared to one-way text outputs [zhang2025human]. Their analysis shows that many existing AI systems remain static and fail to leverage interactive elements that foster user engagement and a deeper understanding of the AI reasoning process. In educational settings, a study demonstrates that interactive LLM-driven question–answer systems increase learners’ focus and cognitive engagement without introducing cognitive load [ozdel2025examining]. Together, these studies highlight that interactivity in AI explanations not only improves the user’s understanding but also keeps the users engaged.
@@
Our work extends these insights by introducing interactive reasoning formats that move beyond passive text outputs. We explore how interactivity can make LLM reasoning both more interpretable and engaging by allowing users to navigate reasoning steps, inspect structure, and verify logic.
3 Method
3.1 Designing Alternative Response Formats for LLM Reasoning
@@
The goal of this project is to design and evaluate alternative response formats for LLM-generated reasoning for math problems. While LLMs have demonstrated strong capabilities in producing step-by-step reasoning through CoT prompting, prior research shows that these outputs often are lengthy, text-heavy passages that are difficult for students to parse and verify in practice [plevris2023chatbots]. Therefore, we investigate whether alternative reasoning formats can enhance comprehension and reduce the cognitive load of users. To address these challenges, we design three alternative interactive reasoning formats that transform the same underlying content into different modes of presentation. We now introduce our three alternative explanation formats alongside the baseline explanation format.
@@
Traditional Chain-of-Thought (CoT) serves as the baseline explanation format. It presents reasoning in a purely text-based form, where the model expresses its thought process step-by-step in natural language. Each step builds upon the previous one and forms a continuous logic flow that leads to the final solution[wei2022chain]. The interface includes three main sections: the problem statement, the model’s step-by-step explanation, and the final answer (see Fig. 2 A). This format represents the standard way large language models naturally generate explanations without additional structural guidance or interactivity.
Interactive Chain-of-Thought (iCoT) demonstrates a more traditional format but restructures the reasoning into discrete, highlighted segments (see Fig. 2 B). Each step is visually separated and includes embedded formulas and calculations, which are revealed sequentially using playback controls. This interface balances the linguistic reasoning with the structured presentation by avoiding long, uninterrupted paragraphs. Users can focus on one step at a time and interactively navigate the solution flow, which makes it easier to detect inconsistencies or mistakes.
@@
Figure 2: Examples of four explanation formats used in the study: (A) traditional Chain-of-Thought (CoT), (B) interactive Chain-of-Thought (iCoT), (C) interactive Program-of-Thought (iPoT), and (D) interactive Graph (iGraph). Each format presents the same reasoning steps in a different modality (textual, structured, code-like, or visual). For consistency, all four formats present the same mathematical problem. Complete interfaces for these four example explanation formats are displayed in Appendix A (Fig. 9,12).
@@
Interactive Program-of-Thought (iPoT) shows explanations in a structured, code-like format(see Fig. 2 C). Each reasoning step is represented as a line of pseudo-code, where variables are defined, updated, and computed systematically. This format allows users to trace how each intermediate value is derived through reasoning steps. Playback controls enable users to execute the reasoning “step-by-step,” similar to debugging or running a script. This interface design draws inspiration from the “Program of Thoughts Prompting” framework proposed by chen2022program, which demonstrated that representing reasoning as executable programs can improve interpretability and consistency in LLMs.
@@ 3.2 Dataset Construction
To evaluate how alternative explanation formats aid users in detecting LLM-generated errors, we construct a dataset derived from GSM8K, a widely used benchmark of grade school math word problems [cobbe2021training]. GSM8K is chosen because it contains a large and diverse set of multi-step reasoning tasks that resemble problems students might encounter in classroom and tutoring settings. Each GSM8K problem typically involves 3–8 steps of reasoning with natural language explanations, which makes it well-suited for transformation into interactive formats.
@@
A central aspect of our design is the controlled injection of errors into the explanations. Rather than relying on naturally occurring model mistakes, following kamoi2024evaluating, we adopt a taxonomy of nine error categories: Calculation (CA), Counting (CO), Context Value (CV), Contradictory Step (CS), Missing Step (MS), Hallucination (HA), Unit Conversion (UC), Operator (OP), and Formula Confusion (FC). By systematically injecting each of these error types, we creat a consistent and fine-grained framework that enable us to assess not only whether participants could recognize when an explanation is incorrect, but also whether certain categories of errors are more easily detected across different interface formats. For each error type, we randomly sample 50 problems from GSM8K. We ensure that each selected problem contains between four to seven reasoning steps. This range is chosen because shorter problems may make error detection trivial, while longer problems risk overburdening participants. Altogether, this process produced 450 erroneous explanations (9 error types × 50 examples each). To balance the dataset and establish a baseline for comparison, we also sample an additional 50 correct explanations from GSM8K. Like the erroneous cases, these problems also contain 4–7 reasoning steps in length and served to evaluate whether participants could correctly recognize valid reasoning when presented in different formats. By combining erroneous and correct cases, we obtain a dataset of 500 explanations in total.
3.3 Tagging as an Intermediate Representation

### v4

@@ Abstract.
The reasoning capabilities of Large Language Models (LLMs) have led to their increasing employment in several critical applications, particularly education, where they support problem-solving, tutoring, and personalized study. While there are a plethora of works showing the effectiveness of LLMs in generating step-by-step solutions through chain-of-thought (CoT) reasoning on reasoning benchmarks, little is understood about whether the generated CoT is helpful for end-users in improving their ability to comprehend mathematical reasoning problems and detect errors/hallucinations in LLM-generated solutions. To address this gap and contribute to understanding how reasoning can improve human-AI interaction, we present three new interactive reasoning interfaces: interactive CoT (iCoT), interactive Program-of-Thought (iPoT), and interactive Graph (iGraph), and a novel framework that generates the LLM’s reasoning from traditional CoT to alternative, interactive formats. Across 125 participants, we found that interactive interfaces significantly improved performance. Specifically, the iGraph interface yielded the highest clarity and error detection rate ($85.6\%$), followed by iPoT ($82.5\%$), iCoT ($80.6\%$), all outperforming standard CoT ($73.5\%$). Interactive interfaces also led to faster response times, where participants using iGraph were fastest ($57.9$ secs), compared to iCoT and iPoT ($60$ secs), and the standard CoT baseline ($64.7$ secs). Furthermore, participants preferred the iGraph reasoning interface, citing its superior ability to enable users to follow the LLM’s reasoning process. We discuss the implications of these results and provide recommendations for the future design of reasoning models. The code and interfaces for this project can be found here: https://github.com/Runtaozhou/Interactive-CoT
Interactive Chain-of-Thought, Interactive Reasoning Interfaces, Explainable Artificial Intelligence (XAI), Human-AI Interaction, Large Language Models (LLMs), Educational AI, Mathematical Reasoning
journalyear: 2026copyright: ccconference: 31st International Conference on Intelligent User Interfaces; March 23–26, 2026; Paphos, Cyprusbooktitle: 31st International Conference on Intelligent User Interfaces (IUI ’26), March 23–26, 2026, Paphos, Cyprusdoi: 10.1145/3742413.3789134isbn: 979-8-4007-1984-4/2026/03ccs: Computing methodologies Artificial intelligenceccs: Human-centered computing Empirical studies in interaction design
1. Introduction
@@
Large language models (LLMs) are now woven into high-stakes, real-world workflows, including education, where they assist with problem solving, tutoring, and personalized study (Kojima et al., 2022; Wang et al., 2024; Forootani, 2025). A key driver of their recent success on complex tasks is chain-of-thought (CoT) prompting (Wang et al., 2023): eliciting step-by-step intermediate reasoning often yields dramatic accuracy gains on arithmetic, commonsense, and symbolic reasoning benchmarks, with state-of-the-art performance on GSM8K and related math datasets. However, despite these advancements, the current format for presenting LLM explanations is far from optimal (Agarwal et al., 2024).
@@
Figure 1. Given a GSM8K question, LLMs typically provide step-by-step reasoning followed by the final answer. However, such output presentation is often static and long, posing a higher cognitive load to users and leading to slower and more erroneous answer verification. In contrast, we prompt LLMs to generate an interactive HTML/JavaScript application wrapped around the reasoning. This interface enables users to verify the reasoning more efficiently via tools of (a) navigation buttons (inspired by those in common IDEs) and (b) colored highlights (Nguyen et al., 2025).
@@
Despite these advances, it remains unclear whether text-only CoT explanations actually help end users—students, teachers, or non-expert verifiers—understand and check mathematical reasoning. While these outputs contain rich information, their format introduces significant challenges for human interpretation (Kunz and Kuhlmann, 2024; Hedlin et al., 2025; Turpin et al., 2023). In practice, CoT often appears as a long “wall of text” that can be time-consuming, obscure structure, increase cognitive load, and make error detection harder precisely when it matters. For example, when an explanation contains a subtle mistake in one intermediate step, detecting that error often requires carefully parsing multiple sentences and mentally reconstructing the reasoning flow (Mishra et al., 2025). More broadly, the explainable AI (XAI) literature shows that explanation techniques are frequently designed around what researchers can produce rather than what end users need; users instead report wanting practically useful information that improves collaboration, trust calibration, and feedback loops with AI systems(Miller, 2019; Abdul et al., 2018; Kaur et al., 2020). This gap motivates rethinking how reasoning is surfaced to end-users in contexts like education, where clarity, verification, and engagement are essential.
@@
Prior work has demonstrated that the design of LLM explanations, which includes factors such as visual structure, interactivity, and information hierarchy, can significantly influence users’ trust and understanding (Yang et al., 2020; Alicioglu and Sun, 2022; Jiang et al., 2023; Nguyen et al., 2024a). However, most existing research on CoT focuses on improving the reasoning process within the model rather than optimizing the presentation of the reasoning (Wang and Zhou, 2024; Wang et al., 2022). As a result, LLM explanations remain largely static, text-heavy, and difficult to navigate (Zhao et al., 2024). This disconnect raises critical questions: How can explanation formats be redesigned to reduce cognitive load and enable users to effectively detect errors and understand LLM-generated reasoning steps?
@@
RQ1) Do interactive explanation formats improve verification accuracy compared to traditional text outputs?
RQ2) How do these formats influence the time and effort required to verify reasoning?
@@ 2. Related Work
LLMs in Education with Chain-of-Thought (CoT) Prompting. A key prompting technique for LLMs is CoT prompting, which elicits the LLM’s reasoning steps and improves its performance on multi-step tasks (Zhang et al., 2022). A recent study shows that few-shot CoT dramatically boosts arithmetic, commonsense, and symbolic reasoning accuracy (Wei et al., 2022; Kojima et al., 2022). In educational settings, however, how explanations are presented matters. Despite clear opportunities for tutoring and feedback, Khosravi et al. (2022) highlights persistent concerns about transparency, interpretability, and over-reliance in LLMs. Beyond transparency, CoT reasoning tends to be quite lengthy, which can increase cognitive load and slow down the user’s comprehension (Nayab et al., 2024). Given the concerns about the interpretability and cognitive demands of LLM-generated reasoning, researchers have begun to explore how reasoning quality and user understanding can be evaluated more systematically. A growing body of work has adopted verification accuracy, which indicates the users’ ability to correctly judge whether a model’s reasoning is valid, as a key indicator of comprehension in LLM explanations (Narayanan et al., 2018; Linder et al., 2021; Nourani et al., 2020; Lage et al., 2019; Nguyen et al., 2024a). Further, a recent study argues that explanations are only useful when they enable users to verify predictions step-by-step, rather than simply accepting plausible reasoning at face value (Fok and Weld, 2024).In our study, we aim to address this question by measuring how effectively people can detect reasoning errors across different explanation formats, allowing us to assess not only the quality of the reasoning itself but also how various interface designs influence users’ ability to verify and understand it.
@@
Interactivity in LLM Explanations. Recent studies have emphasized the importance of interactivity in AI explanations to improve deeper user engagement and understanding(Kim et al., 2020; Guesmi et al., 2024; Nguyen et al., 2024b)Teso et al. (2023) provides a comprehensive overview of interactive machine learning, which shows that explanations are most effective when users can query, edit, or refine them rather than passively receive them. They argue that interaction transforms explanations from static outputs into dynamic learning tools that encourage users’ cognitive engagement. Similarly, another work reviews human-centered explainability and finds that interactive explanation interfaces significantly enhance engagement, comprehension, and retention compared to one-way text outputs (Zhang et al., 2025). Their analysis shows that many existing AI systems remain static and fail to leverage interactive elements that foster user engagement and a deeper understanding of the AI reasoning process. In educational settings, a study demonstrates that interactive LLM-driven question–answer systems increase learners’ focus and cognitive engagement without introducing cognitive load (Ozdel et al., 2025). Together, these studies highlight that interactivity in AI explanations not only improves the user’s understanding but also keeps the users engaged.
Transparency and reliance in LLM Explanations. Prior work in human-centered XAI highlights that transparency and reliance are closely related but not equivalent(de Souza Filho et al., 2024). Recent work claims that transparency emerges from the combined design of models, data, and interfaces, and advocates for simplified and interactive explanations to support user understanding(Muralidhar et al., 2023). Another work argues that explanations can unintentionally induce over- or under-reliance, and that current evidence on trust calibration remains inconclusive without explicit reliance-aware UI design(de Souza Filho et al., 2024). Building on top of these insights, a recent work on selective transparency via progressive disclosure demonstrates that layered explanations reduce cognitive overload while supporting appropriate reliance in clinical settings(Muralidhar et al., 2025).These works suggest LLM explanations must be transparent, context-aware, and designed to calibrate reliance rather than maximize information disclosure.
@@
Our work extends these insights by introducing interactive reasoning formats that move beyond passive text outputs. We explore how interactivity can make LLM reasoning more transparent, interpretable, and engaging by allowing users to navigate reasoning steps, inspect structure, and verify logic.
3. Method
3.1. Designing Alternative Response Formats for LLM Reasoning
@@
The goal of this project is to design and evaluate alternative response formats for LLM-generated reasoning for math problems. While LLMs have demonstrated strong capabilities in producing step-by-step reasoning through CoT prompting, prior research shows that these outputs often are lengthy, text-heavy passages that are difficult for students to parse and verify in practice (Plevris et al., 2023). Therefore, we investigate whether alternative reasoning formats can enhance comprehension and reduce the cognitive load of users. To address these challenges, we design three alternative interactive reasoning formats that transform the same underlying content into different modes of presentation. We now introduce our three alternative explanation formats alongside the baseline explanation format.
@@
Traditional Chain-of-Thought (CoT) serves as the baseline explanation format. It presents reasoning in a purely text-based form, where the model expresses its thought process step-by-step in natural language. Each step builds upon the previous one and forms a continuous logic flow that leads to the final solution(Wei et al., 2022). The interface includes three main sections: the problem statement, the model’s step-by-step explanation, and the final answer (see Fig. 2 A). This format represents the standard way large language models naturally generate explanations without additional structural guidance or interactivity.
Interactive Chain-of-Thought (iCoT) demonstrates a more traditional format but restructures the reasoning into discrete, highlighted segments (see Fig. 2 B). iCoT introduces three key interactive features: (1) a dual-panel design where the left panel summarizes key information and statistics from the question for easy reference; (2) color-coded highlights that emphasize critical information and numbers across both panels, helping users identify and cross-reference key details; and (3) navigation buttons at the top of the interface that display one reasoning step at a time, which allow users to progress through the explanation at their own pace.These features work together with the aim of making it easier for users to detect inconsistencies or mistakes.
@@
Figure 2. Examples of four explanation formats used in the study: (A) traditional Chain-of-Thought (CoT), (B) interactive Chain-of-Thought (iCoT), (C) interactive Program-of-Thought (iPoT), and (D) interactive Graph (iGraph). Each format presents the same reasoning steps in a different modality (textual, structured, code-like, or visual). For consistency, all four formats present the same mathematical problem. Complete interfaces for these four example explanation formats are displayed in appendix A (figs. 9 and 12).
@@
Interactive Program-of-Thought (iPoT) shows explanations in a structured, code-like format (see Fig. 2 C). Each reasoning step is represented as a line of pseudo-code, where variables are defined, updated, and computed systematically. This format allows users to trace how each intermediate value is derived through reasoning steps. Playback controls enable users to execute the reasoning “step-by-step,” similar to debugging or running a script. This interface design draws inspiration from the “Program of Thoughts Prompting” framework proposed by Chen et al. (2022), which demonstrated that representing reasoning as executable programs can improve interpretability and consistency in LLMs.
@@ 3.2. Dataset Construction
To evaluate how alternative explanation formats aid users in detecting LLM-generated errors, we construct a dataset derived from GSM8K, a widely used benchmark of grade school math word problems (Cobbe et al., 2021). GSM8K is chosen because it contains a large and diverse set of multi-step reasoning tasks that resemble problems students might encounter in classroom and tutoring settings. Each GSM8K problem typically involves 3–8 steps of reasoning with natural language explanations, which makes it well-suited for transformation into interactive formats.
@@
A central aspect of our design is the controlled injection of errors into the explanations. Rather than relying on naturally occurring model mistakes, following Li et al. (2024), we adopt the dataset that they create with the taxonomy of nine error categories: Calculation (CA), Counting (CO), Context Value (CV), Contradictory Step (CS), Missing Step (MS), Hallucination (HA), Unit Conversion (UC), Operator (OP), and Formula Confusion (FC). In this dataset, the correct and erroneous reasoning steps have already been labeled and validated. By systematically using each of these error types, we create a consistent and fine-grained framework that enable us to assess not only whether participants could recognize when an explanation is incorrect, but also whether certain categories of errors are more easily detected across different interface formats  [^1] A detailed analysis of verification accuracy for each individual error category is presented in Appendix. C. For each error type, we randomly sample 50 problems from GSM8K. To ensure balanced difficulty across conditions, we sampled explanations containing relatively similar numbers of reasoning steps (between four to seven steps) within each error category. This sampling strategy ensures that participants encounter comparable task complexity regardless of their assigned interface format. This range is chosen because shorter problems may make error detection trivial, while longer problems risk overburdening participants. Altogether, this process produced 450 erroneous explanations (9 error types × 50 examples each). To balance the dataset and establish a baseline for comparison, we also sample an additional 50 correct explanations from GSM8K. Like the erroneous cases, these problems also contain 4–7 reasoning steps in length and served to evaluate whether participants could correctly recognize valid reasoning when presented in different formats. By combining erroneous and correct cases, we obtain a dataset of 500 explanations in total.
3.3. Tagging as an Intermediate Representation